OpenStudy (anonymous):

wrtie an equation of the line parallel to y=3x-1 and passing through the point (2,-5)

OpenStudy (anonymous):

show the general formula please

OpenStudy (anonymous):

parallel = same slope = in this case 3; y = 3 x + b, passes throught the point (2,-5), -5 = 3 * 2 + b -5 = 6 + b b = - 11 y = 3x - 11

OpenStudy (anonymous):

wow ok thank you

OpenStudy (anonymous):

You're welcome.